In the following experiment, a coffee-cup calorimeter containing 100 g of H2O is used. The initial temperature of the calorimeter is 23.0 oC. If 7.6 g of CaCl2is added to the calorimeter, what will be the final temperature of the solution in the calorimeter? Assume that the heat capacity of the solution is 4.18 J/goC, and that the heat capacity of the calorimeter is negligible. The heat of dissolution ΔHsoln of CaCl2 is −82.8 kJ/mol. Enter your answer numerically in terms of oC.
Answer #1

ΔHsoln of CaCl2 = −82.8 kJ/mol.

moles of CaCl2 = 7.6 / 110.98 = 0.0685 mol

ΔHsoln = - Q / n

- 82.8 = - Q / 0.0685

Q = 5.67 kJ

Q = 5670 J

Q = m Cp dT

5670 = 107.6 x 4.18 x dT

dT = 12.6

T2 - 23 = 12.6

T2 = 35.6

final temperature = 35.6 oC